# Cyclicity of Multipliers on the Unit Ball of $\mathbb{C}^{n}$: A Corona-Based Approach

### v2

@@ 2.3. Peak Sets of $A^{\infty}(\mathbb{B}_{n})$
Let $M\subset\partial\mathbb{B}_{n}$ be a smooth real submanifold of real dimension $m$. We say that $M$ is totally real if for every point $p\in M$, the complex structure $J$ satisfies
$T_{p}M\cap J(T_{p}M)=\{0\},$
where $J(v)=iv$ for $v\in\mathbb{C}^{n}$, and $T_{p}M$ is the real tangent space of $M$ at $p$.
@@
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded strictly pseudoconvex domain with smooth boundary $\partial\Omega$. A closed subset $E\subset\partial\Omega$ is called a peak set for the algebra $A^{\infty}(\Omega)$ if there exists a function $f\in A^{\infty}(\Omega)$ such that
@@
$|f(z)|<1\quad\text{for all }z\in\overline{\Omega}\setminus E.$
Equivalently, $E$ is a peak set if there exists a function $g\in A^{\infty}(\Omega)$ such that
@@
$\operatorname{Re}g(z)>0\quad\text{for all }z\in\overline{\Omega}\setminus E.$
@@
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded strictly pseudoconvex domain with $C^{\infty}$ boundary. Let $M\subset\partial\Omega$ be a $C^{\infty}$ submanifold of real dimension at most n1n-1 such that for every point $p\in M$, the tangent space $T_{p}(M)\subset T^{\mathbb{C}}_{p}(\partial\Omega)$. Then every compact subset $K\subset M$ is a peak set for the algebra $A^{\infty}(\Omega)$.
Moreover, if $g$ is the corresponding peak function in the sense of Definition 5, then there exists a constant $c>0$ and an open neighborhood $U\subset\mathbb{C}^{n}$ of $K$ such that
(4) $\operatorname{Re}g(z)\geq c\cdot\operatorname{dist}(z,M)^{2}\quad\text{for all }z\in\Omega\cap U.$
@@
Under the hypotheses of Theorem 9, for every point $z_{0}\in K$, there exist constants $c>0$ and $c^{\prime}>0$, and an open neighborhood $U\subset\mathbb{C}^{n}$ of $z_{0}$, such that for all $z\in U\cap\Omega$, we have
@@
where $w=u+iv$ are local holomorphic coordinates centered at $z_{0}$, with $u$ in the inward normal direction to $\partial\Omega$, and $v$ in the tangential direction.
@@ 3. Proof of the Main Result
Let $g=\sum_{k\geq 0}g_{k}$ where $g_{k}$’s are $k$-homogeneous polynomials. So if we define $\phi=g_{r}Rg-gRg_{r}$, we may write:
$|R^{t}\phi|=|R^{t}((g_{r}-g)Rg+g(Rg-Rg_{r}))|=|R^{t}(Rg\sum_{k\geq 0}(r^{k}-1)g_{k}+g\sum_{k\geq 0}(1-r^{k})kg_{k})|\leq C_{t}(1-r)$
where $C_{t}$ depends only on the order of differentiation $t$. Noting that $|R^{l}\dfrac{g}{g_{r}}|=|R^{l-1}\dfrac{\phi}{g_{r}^{2}}|$ finishes the proof of the lemma since $g$ and all of its derivatives are bounded. ∎
@@
Let $M\subset\partial\mathbb{B}_{n}$ be as in the Theorem 9 with $m\coloneqq\dim_{\mathbb{R}}M\leq n-1$ and $K\subseteq M$ compact then there exists a function $g\in A^{\infty}(\mathbb{B}_{n})$ such that $\mathcal{Z}(g)\cap\overline{\mathbb{B}}_{n}=K$ which is cyclic in $D_{\alpha}(\mathbb{B}_{n})$ for any $\alpha\leq\alpha_{c}=\frac{2n-m}{2}$. In fact, the function $g$ here is the same as the one in Theorem 9.
@@
Corollary 10 ensures the existence of a function $g\in A^{\infty}(\mathbb{B}_{n})$ with the desired zero set. We will show that this function is cyclic in $D_{\alpha_{c}}(\mathbb{B}_{n})$.
@@
(6) $\left\|\frac{g}{g_{r}}\right\|_{\alpha_{c}}\asymp\int_{\mathbb{B}_{n}}(1-|z|^{2})^{2l-\alpha_{c}-1}\left|R^{l}\left(\frac{g}{g_{r}}\right)\right|^{2}\,dv(z).$
@@
(7) $\left\|\frac{g}{g_{r}}\right\|_{\alpha_{c}}\lesssim(1-r)^{2}\int_{\mathbb{B}_{n}}(1-|z|^{2})^{2l-\alpha_{c}-1}\left|\frac{1}{g_{r}}\right|^{2l+2}\,dv(z).$
Now, for $p\in M$, let $Z(x,y,w)$ be a local coordinate at $p$, with $x\in\mathbb{R}^{m}$, $y\in\mathbb{R}^{2n-m-2}$, and $w=u+iv\in\mathbb{C}$, such that $Z(x,0,0)\in M$, $Z(x,y,0)\in\partial\mathbb{B}_{n}$, and $w$ is as in Corollary 10. Then, for $r>\frac{1}{2}$, there exists $c>0$ such that
(8) $|g_{r}(Z)|\geq c\left((1-r)+u+|v|+|y|^{2}\right).$
To estimate the integral in (6), it suffices to restrict it to the region where the local coordinate $Z$ is valid. Thus, we write:
(9) $\left\|\frac{g}{g_{r}}\right\|_{\alpha_{c}}\lesssim(1-r)^{2}\int_{(0,1)^{m}}\int_{(0,1)^{2n-m-2}}\int_{(-1,1)}\int_{(0,1)}\frac{u^{2l-\alpha_{c}-1}}{((1-r)+u+|v|+|y|^{2})^{2l+2}}\,du\,dv\,dy\,dx.$
@@
$\left\|\frac{g}{g_{r}}\right\|_{\alpha_{c}}$ $\lesssim(1-r)^{2}\int_{(0,1)^{m}}\int_{(-1,1)^{2n-m-2}}\int_{(-1,1)}\int_{(0,1)}\frac{1}{((1-r)+u^{1/\gamma}+|v|+|y|^{2})^{2l+2}}\,du\,dv\,dy\,dx$
@@
$\asymp(1-r)^{2}\int_{(0,1)^{m}}\int_{(-1,1)^{2n-m-2}}\frac{1}{((1-r)+y_{1}^{2}+\cdots+y_{2n-m-2}^{2})^{\alpha_{c}+1}}\,dy\,dx$
@@
By Theorem 9, we know that $\mathcal{Z}(f)\cap\overline{\mathbb{B}}_{n}=M$, is a peak set by Theorem 9. Therefore, by Lemma 12, there exists $g\in A^{\infty}(\mathbb{B}_{n})$ such that $g$ is cyclic in $D_{\frac{2n-m}{2}}(\mathbb{B}_{n})$, $\mathcal{Z}(g)\cap\overline{\mathbb{B}}_{n}=M$, and $g$ is a cyclic multiplier.
Note that for $g$, by construction, we have:
@@
Therefore, possibly after raising $g$ to a suitable power and multiplying by a positive constant, we have on $\mathbb{B}_{n}$:
$|g(z)|\leq|f(z)|.$
Noting that $f$ and $g$ satisfy the conditions of Theorem 3, and that $g$ is cyclic (hence so is $g^{N}$, since $g$ is a multiplier), completes the proof. ∎
@@ Acknowledgements
The author would like to express sincere gratitude to Professors Łukasz Kosiński and Pascal Thomas, the author’s PhD advisors, for their invaluable guidance and ongoing support throughout this work. The author is particularly grateful to Professor Thomas for his substantial help in the development of the present idea, and to Professor Kosiński for his continued insightful advice and encouragement. Finally, the author acknowledges the Institut de Mathématiques de Toulouse and the Department of Mathematics at Jagiellonian University for providing a stimulating research environment.

### v3

@@ 2.3. Peak Sets of $A^{\infty}(\mathbb{B}_{n})$
For $\zeta\in\partial\mathbb{B}_{n}$, the complex tangent space to the sphere is defined by
$T_{\zeta}^{\mathbb{C}}\partial\mathbb{B}_{n}:=\{v\in\mathbb{C}^{n}:\langle v,\zeta\rangle=0\}.$
@@
Let $M\subset\mathbb{\partial}\mathbb{B}_{n}$ be a $\mathcal{C}^{k}$ real submanifold. We say that $M$ is complex tangential if, for every $\zeta\in M$,
$T_{\zeta}M\subset T_{\zeta}^{\mathbb{C}}\partial\mathbb{B}_{n}.$
Definition 20.
Let $M\subset\mathbb{\partial}\mathbb{B}_{n}$ be a $\mathcal{C}^{k}$ real submanifold. We say that $M$ is totally real if, for every $\zeta\in M$,
$T_{\zeta}M\cap iT_{\zeta}M=\{0\}.$
Definition 21.
A closed subset $E\subset\partial\mathbb{B}_{n}$ is called a peak set for $A^{\infty}(\mathbb{B}_{n})$ if there exists a function $f\in A^{\infty}(\mathbb{B}_{n})$ such that
@@
$|f(z)|<1\quad\text{for all }z\in\overline{\mathbb{B}}_{n}\setminus E.$
Equivalently, $E$ is a peak set if there exists a function $g\in A^{\infty}(\mathbb{B}_{n})$ such that
@@
$\operatorname{Re}g(z)>0\quad\text{for all }z\in\overline{\mathbb{B}}_{n}\setminus E.$
@@
Let $M\subset\partial\mathbb{B}_{n}$ be a $C^{\infty}$ complex tangential manifold of real dimension at most n1n-1. Then every compact subset $K\subset M$ is a peak set for $A^{\infty}(\mathbb{B}_{n})$.
Moreover, it follows from equation (21.4) in [6], together with the discussion on page 182 therein, that for a corresponding peak function $g$ in the sense of Definition 21, there exist a constant $c>0$ and an open neighborhood $U\subset\mathbb{C}^{n}$ of $K$ such that
(5) $\operatorname{Re}g(z)\geq c\cdot\operatorname{dist}(z,M)^{2}\quad\text{for all }z\in\mathbb{B}_{n}\cap U.$
@@
Under the hypotheses of Theorem 22, for every point $z_{0}\in K$, there exist constants $c>0$ and $c^{\prime}>0$, and an open neighborhood $U\subset\mathbb{C}^{n}$ of $z_{0}$, such that for all $z\in U\cap\mathbb{B}_{n}$, we have
@@
where $w=u+iv$ are local holomorphic coordinates centered at $z_{0}$, with $u$ in the inward normal direction to $\partial\mathbb{B}_{n}$, and $v$ in the tangential direction and transverse to $M$ by the hypothesis on $M$.
@@ 3. Proof of the Main Result
Let $\phi=g_{r}Rg-gRg_{r}$. For an integer $0\leq t\leq l$, we may write:
(7) $|R^{t}\phi|$ $=\left|R^{t}\left((g_{r}-g)Rg+g(Rg-Rg_{r})\right)\right|$
$\leq\left|R^{t}\left((g_{r}-g)Rg\right)\right|+\left|R^{t}\left(g(Rg-Rg_{r})\right)\right|$
$\leq\sum_{s=0}^{t}\left|R^{s}(g-g_{r})\right|\left|R^{t+1-s}g\right|+\sum_{s=0}^{t}\left|R^{s}(Rg-Rg_{r})\right|\left|R^{t-s}g\right|.$
Since $g$ and all its derivatives are bounded, there exists $C>0$ such that
(8) $|R^{s}g|(z)\leq C,$
for all $z\in\overline{\mathbb{B}}_{n}$ and all $s\in\{1,\ldots,l+1\}$. Also, for $s\in\{1,\ldots,l+1\}$,
(9) $|R^{s}(g-g_{r})(z)|=|R^{s}g(z)-R^{s}g(rz)|\leq\sup_{w\in\mathbb{B}_{n}}|\nabla(R^{s}g)(w)|\,|z-rz|\leq C^{\prime}(1-r),$
where $C^{\prime}>0$ depends only on $l$ and $g$. From (7), (8), and (9), we infer that
$|R^{t}\phi|\leq C^{\prime}_{l}(1-r).$
Noting that $|R^{l}\dfrac{g}{g_{r}}|=|R^{l-1}\dfrac{\phi}{g_{r}^{2}}|$ finishes the proof. ∎
@@
Let $M\subset\partial\mathbb{B}_{n}$, $K\subseteq M$, and $g$ be as in Theorem 22, with $m\coloneqq\dim_{\mathbb{R}}M\leq n-1$. Then $\mathcal{Z}(g)\cap\overline{\mathbb{B}}_{n}=K$, and $g$ is cyclic in $D_{\alpha}(\mathbb{B}_{n})$ for any $\alpha\leq\alpha_{c}=\frac{2n-m}{2}$.
@@
The fact that $\mathcal{Z}(g)\cap\overline{\mathbb{B}}_{n}=K$ comes directly from the fact that $g$ is a peak function for the set $K$. We will show that this function is cyclic in $D_{\alpha_{c}}(\mathbb{B}_{n})$. Note that this proves the cyclicity of $g$ in $D_{\alpha}(\mathbb{B}_{n})$ for any $\alpha\leq\alpha_{c}$. Indeed, this follows from the fact that in this case $\|\cdot\|_{\alpha}\leq\|\cdot\|_{\alpha_{c}}$. Hence, if there exists a sequence of polynomials $(p_{n})$ such that $\|1-p_{n}g\|_{\alpha_{c}}\to 0$, then we also have $\|1-p_{n}g\|_{\alpha}\to 0$.
@@
(10) $\left\|\frac{g}{g_{r}}\right\|^{2}_{\alpha_{c}}\asymp\int_{\mathbb{B}_{n}}(1-|z|^{2})^{2l-\alpha_{c}-1}\left|R^{l}\left(\frac{g}{g_{r}}\right)\right|^{2}\,dv(z).$
@@
(11) $\left\|\frac{g}{g_{r}}\right\|^{2}_{\alpha_{c}}\lesssim(1-r)^{2}\int_{\mathbb{B}_{n}}(1-|z|^{2})^{2l-\alpha_{c}-1}\left|\frac{1}{g_{r}}\right|^{2l+2}\,dv(z).$
Now, for $p\in M$, let $Z_{p}(x,y,w)$ be a system of local coordinates at $p$, defined on an open neighborhood $U_{p}$ of $p$, where $x\in\mathbb{R}^{m}$, $y\in\mathbb{R}^{2n-m-2}$, and $w=u+iv\in\mathbb{C}$, such that
$Z_{p}(x,0,0)\in M,\qquad Z_{p}(x,y,0)\in\partial\mathbb{B}_{n},$
and where $w$ is as in Corollary 23. We may choose $0<r_{p}<1$ sufficiently large such that, for any $r_{p}\leq r<1$, we have $rZ_{p}\in U_{p}$. Then there exists $c_{p}>0$ such that
(12) $|g_{r}(Z_{p})|\geq c_{p}\left((1-r)+u+|v|+|y|^{2}\right).$
Since $M$ is compact, we may cover it by finitely many such neighborhoods $U_{p}$. This allows us to estimate the integral in (10) in the following form:
(13) $\left\|\frac{g}{g_{r}}\right\|^{2}_{\alpha_{c}}\lesssim(1-r)^{2}\int_{(0,1)^{m}}\int_{(-1,1)^{2n-m-2}}\int_{(-1,1)}\int_{(0,1)}\frac{u^{2l-\alpha_{c}-1}}{((1-r)+u+|v|+|y|^{2})^{2l+2}}\,du\,dv\,dy\,dx.$
@@
$\left\|\frac{g}{g_{r}}\right\|^{2}_{\alpha_{c}}$ $\lesssim(1-r)^{2}\int_{(0,1)^{m}}\int_{(-1,1)^{2n-m-2}}\int_{(-1,1)}\int_{(0,1)}\frac{1}{((1-r)+u^{1/\gamma}+|v|+|y|^{2})^{2l+2}}\,du\,dv\,dy\,dx$
@@
$=(1-r)^{2}\int_{(0,1)^{m}}\int_{(-1,1)^{2n-m-2}}\frac{1}{((1-r)+y_{1}^{2}+\cdots+y_{2n-m-2}^{2})^{\alpha_{c}+1}}\,dy\,dx$
@@
From Theorem 22, we know that $\mathcal{Z}(f)\cap\overline{\mathbb{B}}_{n}=M$ is a peak set. Therefore, by Lemma 25, there exists $g\in A^{\infty}(\mathbb{B}_{n})$ such that $\mathcal{Z}(g)\cap\overline{\mathbb{B}}_{n}=M$, and $g$ is a cyclic multiplier of $D_{\frac{2n-m}{2}}(\mathbb{B}_{n})$.
From the construction of $g$, we obtain:
@@
Therefore, there exists $c>0$ such that we have on $\mathbb{B}_{n}$:
$c|g^{l}(z)|\leq|f(z)|.$
Replacing $g$ by $cg^{l}$ and keeping the notation $g$, we may assume that $f$ and $g$ satisfy the assumptions of Theorem 3. Hence, there exists $N\in\mathbb{N}$ such that
$g^{N}]\subseteq[f].$
Finally, since $g$ is cyclic (and therefore so is $g^{N}$, as $g$ is a multiplier), the proof is complete. ∎
@@ Acknowledgements
The author would like to express sincere gratitude to Łukasz Kosiński and Pascal Thomas, the author’s PhD advisors, for their invaluable guidance and ongoing support throughout this work. The author is particularly grateful to Pascal Thomas for his substantial help in the development of the present idea, and to Łukasz Kosiński for his continued insightful advice and encouragement. Finally, the author acknowledges the Institut de Mathématiques de Toulouse and the Department of Mathematics at Jagiellonian University for providing a stimulating research environment.